How would I solve: 2 ^3x-5 = 7
Answer by stanbon(75887)

2^(3x-5) = 7
------
Take the long of both sides::
(3x-5)*log(2) = log(7)
----
3x-5 = log(7)/log(2) = 2.807
----
3x = 7.807
-----
x = 2.602
--------------
Cheers,
Stan H.
---------------